# Plastic Arbor: a modern simulation framework for synaptic plasticity – from single synapses to networks of morphological neurons

### v2

@@ Abstract
To enable the modeling of large-scale networks of morphologically detailed neurons with diverse plasticity processes, we have extended the Arbor library to the Plastic Arbor framework, supporting simulations of a large variety of spike-driven plasticity paradigms. To showcase the features of the new framework, we present examples of computational models, beginning with single-synapse dynamics, progressing to multi-synapse rules, and finally scaling up to large recurrent networks. While cross-validating our implementations by comparison with other simulators, we show that Arbor allows simulating plastic networks of multi-compartment neurons at nearly no additional cost in runtime compared to point-neuron simulations. Using the new framework, we have already been able to investigate the impact of dendritic structures on network dynamics across a timescale of several hours, showing a relation between the length of dendritic trees and the ability of the network to efficiently store information.
@@ 1 Introduction
Over the past decades, a number of open-source software simulators have been developed to facilitate the investigation of biological neural networks. Current prominent examples are (Core) NEURON [1, 2], NEST [3], and Brian 2 [4]. While NEST and Brian 2 are widely used for the simulation of large-scale networks of point neurons, NEURON is a well-established tool for modeling neurons with detailed morphology, with its first version already released in the 1980s [1]. Notably, NEURON enables to flexibly create realistic neuron models in its own script language called NMODL. Over the past decades, NEURON was extended to include more features as well as to keep up with the fast development in computing hardware [2, 5]. This has, however, yielded a complex code base that constrains usability, flexibility, and the optimization for modern hardware backends. To overcome these limitations, Arbor has been developed, which is a new simulator library for networks of neurons with detailed morphology. With a Python frontend and support for various model-descriptive formats, including NMODL, Arbor facilitates the implementation and customization of neuron and synapse models. At the same time, Arbor offers heavily optimized execution on different hardware systems [6]. It supports, in particular, modern backend architectures such as multi-core central processing units (CPUs), graphics processing units (GPUs), and message passing interface (MPI) ranks [6], and is openly developed and available on GitHub [7].
@@
Arbor offers a powerful tool for capturing processes at the synaptic, dendritic, and neuronal levels, enabling the examination of their links to network dynamics. Although the core functionality for running network simulations has been available since the inception of the Arbor project [6], the mechanisms required for modeling plasticity-related processes have largely been absent. In our present work, we fill this gap extending the Arbor simulator by the general functionality needed to model diverse spike-driven plasticity rules (e.g., [20, 21, 22]), which constitutes the Plastic Arbor framework. By this, we provide a basis for investigating the functional impact of plasticity dynamics in large networks of morphological neurons via our publicly available code (see the data availability statement below, or visit https://github.com/tetzlab/plastic_arbor).
In this article, we describe the novel functionality in two different sections, where we present methods and results in a side-by-side fashion. The new technical features needed to implement plasticity rules are described in the section ‘Extensions of the Arbor core code’. In the section ‘Computational modeling with synaptic plasticity’, we present examples for the newly implemented technical features introduced before, with increasing complexity of the implemented models. We start from the level of simulating a widely-used spike-timing-dependent plasticity rule in a single synapse, then consider plasticity rules involving multiple synapses, move on to more complex plasticity rules including several hidden states, and finally reproduce findings from large recurrently connected networks. We cross-validate all Arbor implementations with the Brian 2 simulator [4] or with model-specific custom simulators. Importantly, the newly extended functionality of Arbor has enabled us to provide first predictions about the network dynamics underlying synaptic memory consolidation across networks of neurons of different morphology. We specifically provide insights into how the performance in a pattern completion task at the network level depends on the length of the dendrites as well as on the overall size of the employed neurons.
2 Extensions of the Arbor core code
In this section, we present new extensions of the Arbor core code, which are necessary for the implementation of synaptic plasticity models in Arbor. Examples of related models are presented in the following section 3.
@@ 2.4 Diffusion of arbitrary particles
3 Computational modeling with synaptic plasticity
@@
Please note that since plasticity dynamics are the main target of this paper, we only present the model equations that correspond to those. For parts of the considered models other than the plasticity dynamics, please refer to the cited literature and provided code.
@@ 3.1 Spike-timing-dependent plasticity
Spike-timing-dependent plasticity (STDP) is a phenomenon that is reported in a number of experimental studies [32, 33, 34, 23] and described by various theoretical models. The earlier models of STDP have provided a relatively simple form of synaptic plasticity that only depends on the specific timing of pre- and postsynaptic spikes via the temporal differences () between pre- and postsynaptic spikes. In this case, plasticity does not directly depend on the spike rate. As a first step, we implemented a commonly used description [24] given by:
@@
To simulate this model with Arbor, we implemented a single, conductance-based excitatory synapse. This synapse was connected to a Leaky Integrate-and-Fire (LIF) neuron that was stimulated with Poisson input (see Fig. 2a). In addition, an inhibitory synapse was introduced to stabilize the dynamics. Both the excitatory synapse and the LIF neuron were implemented as custom mechanisms in Arbor (defined via NMODL scripts). Note that for this implementation, the POST_EVENT hook, which we added to Arbor’s NMODL dialect, is needed (cf. subsection 2.1) – the hook is called whenever a threshold detector on a cell is triggered. In the present case, this is when the postsynaptic neuron spikes.
@@
In addition to this experiment, we simulated a range of time delay settings to obtain a classical STDP time window, which shows a very good match with Brian 2 simulations as well (Fig. 2d).
@@ 3.2 Spike-driven homeostatic plasticity
After implementing an STDP rule at a single synapse, we now consider a type of synaptic plasticity that depends on multiple presynaptic inputs onto the same postsynaptic neuron.
@@
The resulting weight and firing rate dynamics, along with the varying input rate, are shown in Fig. 2g-i and Supplementary Fig. S1e. We can see that in the case with homeostasis, the resulting firing rate is maintained at values around (Fig. 2i), while in the case without homeostasis (Supplementary Fig. S1e), the resulting firing rate is mainly imposed by the input rate. Note that for the time period with input rate (from to ), the homeostatic weight adjustment can only happen to a limited extent since we do not allow the weights to increase beyond . We cross-validated our Arbor implementation with Brian 2, drawing from an existing example implementation [39].
@@ 3.3 Calcium-based synaptic plasticity
Following the implementation of a simple phenomenological STDP rule (subsection 3.1) as well as spike-driven homeostatic plasticity (subsection 3.2), we now consider a more complex rule of synaptic plasticity that requires multiple postsynaptic variables. This rule describes the potentiation and depression of synaptic strength depending on the postsynaptic calcium concentration, which is driven by pre- and postsynaptic spiking activity. It was presented by Graupner and Brunel in 2012 [20] and has since been used widely. In comparison to the phenomenological STDP rule, the calcium-based rule adapts the synaptic weight dynamics to depend on spike timing and spike rates, both of which have been shown to be important features of long-term synaptic plasticity [40].
@@ 3.6 Synaptic memory consolidation in networks of morphological neurons
Although our present study has been focused on simulation methods, these last findings may provide interesting insights into the role of neuronal structure and dynamics for cognitive functionality at the network level. As the presented results show, Arbor enables to seamlessly move from single- to multi-compartment neurons in a complex network model, leaving the remaining parts of the model unchanged. In the future, the framework that we have developed can be used as a basis for further investigations on neural networks involving diffusion dynamics in multi-compartment neurons.
@@
3.7 Runtime and memory benchmarking
As a final step, we measure the resources required by Arbor to simulate networks of single- and multi-compartment neurons in different computing environments. To this end, we consider the networks of single- or multi-compartment neurons with the 10s-recall paradigm that we used before (subsections 3.5 & 3.6). After running these network models with different Arbor backends on different hardware systems, we compare their runtime and memory use. In the single-compartment case, we also compare to Brian 2 with cpp_standalone device [77, 4] as well as to the custom stand-alone simulator [58] considered before.
Fig. 8 shows that the point-neuron simulators (Brian 2 and stand-alone) need shorter runtimes and less memory compared to Arbor. This is to be expected: Since point-neuron simulators consider neurons without any geometrical structure, they theoretically require much less calculation steps than Arbor, which accounts for at least one finite-size compartment per neuron. Furthermore, the custom stand-alone simulator is highly optimized for the particular model, and can therefore be thought to set an upper bound. Nevertheless, we found that Arbor’s capability of employing GPU hardware can boost its runtime to be even faster than Brian 2 and the custom simulator (Fig. 8a; comparing the main memory use with the GPU case is not meaningful since the GPU has its own memory).
Another point to be mentioned is Arbor’s support for single instruction, multiple data (SIMD) vectorization. As shown in Fig. 8a, switching on SIMD vectorization provides a small but solid improvement in runtime. This comes without any cost for the end user, given that they have a CPU that supports SIMD, which has been the industry standard for many years. The only drawback of SIMD usage in Arbor might be its negative impact on code readability when developing custom mechanism code in C++, which is rather a niche case.
Importantly, our results show that Arbor allows to shift from single-compartment neurons to morphological neurons at almost no cost: both the runtime (Fig. 8a) and the memory consumption (Fig. 8b) only increase slightly when shifting from single-compartment neurons to neurons with compartments. However, it is important to note that the considered multi-compartment network in general exhibits fewer spikes than the single-compartment version (Supplementary Fig. S12b; also cf. subsection 3.6). Thus, given that the number of spikes is a critical factor for the runtime, we checked if the runtime per spike follows as similar trend, and indeed found that this measure also exhibits a slight increase only (Supplementary Fig. S12a).
@@ 3.7 Runtime and memory benchmarking
4 Discussion

### v3

@@ Abstract
To enable the modeling of large-scale networks of morphologically detailed neurons with diverse plasticity processes, we have extended the Arbor library to support simulations of a large variety of spike-driven plasticity paradigms. To showcase the features of the extended framework, we present examples of computational models, beginning with single-synapse dynamics, progressing to multi-synapse rules, and finally scaling up to large recurrent networks. While cross-validating our implementations by comparison with other simulators, we show that Arbor allows simulating plastic networks of multi-compartment neurons at nearly no additional cost in runtime compared to point-neuron simulations. In addition, we demonstrate that Arbor is highly efficient in terms of runtime and memory use as compared to other simulators.
Using the extended framework, as an example, we investigate the impact of dendritic structures on network dynamics across a timescale of several hours, finding a relation between the length of dendritic trees and the ability of the network to efficiently store information.
@@
Author Summary
Understanding how synaptic plasticity shapes the dynamics of the brain requires simulation tools that provide efficient, scalable computation to integrate detailed neuronal morphology into large networks of neurons. Therefore, we extended the open‑source Arbor simulator to enable modeling a comprehensive range of spike‑driven plasticity mechanisms. Using progressively more complex examples – from a single synapse to recurrent networks of thousands of morphological neurons – we demonstrate that the extended Arbor framework matches the results of established simulators. At the same time, it uncloses a regime of new investigations. Benchmarking on modern hardware shows that adding morphology incurs only modest runtime and memory overhead as compared to point-neuron models. Moreover, a comparison with the widely used NEURON framework shows that Arbor is markedly more efficient, even when computing additional plasticity dynamics. Leveraging this performance, as an example, we explore how dendritic length and cell size influence long‑term memory recall, revealing that larger dendritic trees can impair, whereas larger cell diameters can enhance, pattern‑completion performance. The extended Arbor framework and the considered examples of plasticity models are freely available and shall serve to provide a powerful platform for researchers investigating the interplay of cellular structure, synaptic plasticity, and network dynamics.
@@ 1 Introduction
Over the past decades, a number of open-source software simulators have been developed to facilitate the investigation of biological neural networks. Current prominent examples are (Core) NEURON [1, 2], NEST [3], and Brian 2 [4]. While NEST and Brian 2 are widely used for the simulation of large-scale networks of point neurons, NEURON is a well-established tool for modeling neurons with detailed morphology, with its first version already released in the 1980s [1]. Notably, NEURON enables to flexibly create realistic neuron models in its own script language called NMODL. Over the past decades, NEURON was extended to include more features as well as to keep up with the fast development in computing hardware [2, 5]. This has, however, yielded a complex code base that constrains usability, flexibility, and the optimization for modern hardware backends. To overcome these limitations, Arbor has been developed, which is a new simulator library for networks of neurons with detailed morphology. With a Python frontend and support for various model-descriptive formats, including NMODL, Arbor facilitates the implementation and customization of neuron and synapse models. At the same time, Arbor offers heavily optimized execution on different hardware systems [6]. It supports, in particular, modern backend architectures such as multi-core central processing units (CPUs), graphics processing units (GPUs), and message passing interface (MPI) ranks [6], and is openly developed and available on GitHub [7]. Besides CoreNEURON, Arbor is the only simulator with comprehensive GPU and MPI support for multi-compartment models.
@@
Arbor offers a powerful tool for capturing processes at the synaptic, dendritic, and neuronal levels, enabling the examination of their links to network dynamics. Although the core functionality for running network simulations has been available since the inception of the Arbor project [6], the mechanisms required for modeling plasticity-related processes have largely been absent. In our present work, we fill this gap extending the Arbor simulator by the general functionality needed to model diverse spike-driven plasticity rules (e.g., [20, 21, 22]), which constitutes the ‘Plastic Arbor’ framework. By this, we provide a basis for investigating the functional impact of plasticity dynamics in large networks of morphological neurons via our publicly available code (see the data availability statement below, or visit https://github.com/tetzlab/plastic_arbor).
In this article, we aim to present Arbor’s new set of features to model plasticity in large-scale networks of multi-compartment neurons, as well as to demonstrate that Arbor enables highly efficient simulation of such networks in terms of runtime and memory use. We describe the novel functionality in two different sections, where we present methods and results in a side-by-side fashion. The new technical features needed to implement plasticity rules are described in the section ‘Design and Implementation: Extensions of the Arbor core code’. In the section ‘Results: Computational modeling with synaptic plasticity’, we present examples for the newly implemented technical features introduced before, with increasing complexity of the implemented models. We start from the level of simulating a widely-used spike-timing-dependent plasticity rule in a single synapse, then consider plasticity rules involving multiple synapses, move on to more complex plasticity rules including several hidden states, and finally reproduce findings from large recurrently connected networks. We cross-validate all Arbor implementations with the Brian 2 simulator [4] or with model-specific custom simulators. Importantly, the newly extended functionality of Arbor has enabled us to provide first predictions about the network dynamics underlying synaptic memory consolidation across networks of neurons of different morphology. We specifically provide insights into how the performance in a pattern completion task at the network level depends on the length of the dendrites as well as on the overall size of the employed neurons. Finally, we provide benchmarking results showcasing Arbor’s runtime and memory efficiency compared to other simulators.
2 Design and Implementation: Extensions of the Arbor core code
In this section, we present the new extensions of the Arbor core code that are necessary for the implementation of synaptic plasticity models in Arbor. Examples of related models are presented in the following section 3.
@@ 2.4 Diffusion of arbitrary particles
3 Results: Computational modeling with synaptic plasticity
@@
Furthermore, we provide benchmarking results to compare the runtime and memory performance of the simulation of different network models in Arbor and other simulators.
Please note that since the implementation of plasticity dynamics is the main target of this paper, we only present the model equations that correspond to those. For parts of the considered models other than the plasticity dynamics, please refer to the cited literature and provided code.
@@ 3.1 Spike-timing-dependent plasticity
Spike-timing-dependent plasticity (STDP) is a phenomenon that is reported in a number of experimental studies [32, 33, 34, 23] and described by various theoretical models (cf. [35]). The earlier models of STDP have provided a relatively simple form of synaptic plasticity that only depends on the specific timing of pre- and postsynaptic spikes via the temporal differences () between pre- and postsynaptic spikes. In this case, plasticity does not directly depend on the spike rate. As a first step, we implemented a commonly used description [24] given by:
@@
To simulate this model with Arbor, we implemented a single, conductance-based excitatory synapse. This synapse was connected to a Leaky Integrate-and-Fire (LIF) neuron and was stimulated with Poissonian spike input (see Fig. 2a). In addition, an inhibitory synapse was introduced to stabilize the dynamics, also driven by Poissonian spike input. Both the excitatory synapse and the LIF neuron were implemented as custom mechanisms in Arbor (defined via NMODL scripts). Note that for this implementation, the POST_EVENT hook, which we added to Arbor’s NMODL dialect, is needed (cf. subsection 2.1) – the hook is called whenever a threshold detector on a cell is triggered. In the present case, this is when the postsynaptic neuron spikes.
@@
In addition to this experiment, we simulated a range of time delay settings to obtain a classical STDP time window, which also shows a very good match with Brian 2 simulations (Fig. 2d) as well as with the theoretical expectation (Supplementary Fig. S1e).
@@ 3.2 Spike-driven homeostatic plasticity
After implementing an STDP rule at a single synapse, we now consider a type of synaptic plasticity that depends on multiple presynaptic stimuli onto the same postsynaptic neuron.
@@
The resulting weight and firing rate dynamics, along with the varying input rate, are shown in Fig. 2g-i and Supplementary Fig. S1f. We can see that in the case with homeostasis, the resulting firing rate is maintained at values around (Fig. 2i), while in the case without homeostasis (Supplementary Fig. S1f), the resulting firing rate is mainly imposed by the input rate. Note that for the time period with input rate (from to ), the homeostatic weight adjustment can only happen to a limited extent since we do not allow the weights to increase beyond . We cross-validated our Arbor implementation with Brian 2, drawing from an existing example implementation [40].
@@ 3.3 Calcium-based synaptic plasticity
Following the implementation of a simple phenomenological STDP rule (subsection 3.1) as well as spike-driven homeostatic plasticity (subsection 3.2), we now consider a more complex rule of synaptic plasticity that requires multiple postsynaptic variables. This rule describes the potentiation and depression of synaptic strength depending on the postsynaptic calcium concentration, which is driven by pre- and postsynaptic spiking activity. It was presented by Graupner and Brunel in 2012 [20] and has since been used widely. In comparison to the phenomenological STDP rule, the calcium-based rule adapts the synaptic weight dynamics such that they depend on spike timing and spike rates, both of which have been shown to be important features of long-term synaptic plasticity [41].
@@ 3.6 Synaptic memory consolidation in networks of morphological neurons
Although our present study has been focused on simulation methods, these last findings may provide interesting theoretical insights into the role of neuronal structure and dynamics for cognitive functionality at the network level. As the presented results show, Arbor enables to seamlessly move from single- to multi-compartment neurons in a complex network model, leaving the remaining parts of the model unchanged. In the future, the framework that we have developed can be used as a basis for further investigations on neural networks involving diffusion dynamics in multi-compartment neurons.
@@
3.7 Runtime and memory benchmarking with the synaptic memory consolidation model
Employing the network model introduced in the previous section, we scrutinize the resources required by Arbor to simulate networks of single- and multi-compartment neurons in different computing environments. To this end, we consider the networks of single- or multi-compartment neurons with the 10s-recall paradigm that we used before (subsections 3.5 & 3.6). After running these network models with different Arbor backends on different hardware systems, we compare their runtime and memory use. In the single-compartment case, we also compare to Brian 2 with cpp_standalone device [78, 4] as well as to the custom stand-alone simulator [59] considered before.
Fig. 8 shows that the point-neuron simulators (Brian 2 and stand-alone) need shorter runtimes and less memory compared to Arbor. This is to be expected: Since point-neuron simulators consider neurons without any geometrical structure, they theoretically require much less calculation steps than Arbor, which accounts for at least one finite-size compartment per neuron. Furthermore, the custom stand-alone simulator is highly optimized for the particular model, and can therefore be thought to set an upper bound. Nevertheless, we found that Arbor’s capability of employing GPU hardware can boost its runtime to be even faster than Brian 2 and the custom simulator (Fig. 8).
Another point to be mentioned is Arbor’s support for single instruction, multiple data (SIMD) vectorization. As shown in Fig. 8, switching on SIMD vectorization provides a small but solid improvement in runtime. This comes without any cost for the end user, given that they have a CPU that supports SIMD, which has been the industry standard for many years. The only drawback of SIMD usage in Arbor might be its negative impact on code readability when developing custom mechanism code in C++, which is rather a niche case.
Importantly, our results show that Arbor allows to shift from single-compartment neurons to morphological neurons at almost no cost: both the runtime and the memory consumption only increase slightly when shifting from single-compartment neurons (Fig. 8, left-hand side) to neurons with compartments (Fig. 8, right-hand side). However, it is important to note that the considered multi-compartment network in general exhibits fewer spikes than the single-compartment version (Supplementary Fig. S12b; also cf. subsection 3.6). Thus, given that the number of spikes is a critical factor for the runtime, we checked if the runtime per spike follows as similar trend, and indeed found that this measure also exhibits a slight increase only (Supplementary Fig. S12a).
@@ 3.7 Runtime and memory benchmarking with the synaptic memory consolidation model
3.8 Runtime and memory benchmarking with large-scale networks
To underline our claims of Arbor’s scalability and efficiency, in this final section, we consider a set of technical benchmarks based on the so-called busyring benchmark. This is a tunable, well-understood workload that enables stressing various parts of the underlying hardware architecture. It has been used in this capacity as an acceptance benchmark for the novel JUPITER system at the Jülich Supercomputing Centre (JSC) [79].
The basic unit of work in the busyring benchmark is a ring of neurons tuned to propagate a single spike indefinitely (see Fig. 9e). Connections within the ring are realized via conductance-based exponential synapses with a uniform delay and weight . Thus, each neuron spikes at frequency . In addition, a set of connections are placed between ring neurons and random endpoints, where . These connections generate computational load in the code paths responsible for spike transmission and delivery, yet have no effect on the target synapses. The cell model is generated randomly as a tree of given depth, which is done individually per cell to avoid early-stage optimizations. The neuron is then endowed with Hodgkin-Huxley dynamics [80] on the soma region and a plain leak current on the remaining morphology. We call this neuron model simple-branchy. To consider the computational load of plasticity dynamics, we further implemented a variant of the benchmark where the synapses targeted by connections use an STDP model (see subsection 3.1).
| Description | Symbol | Value |
| Number of CPU threads | - | – |
| Number of MPI ranks | - | – |
| Simulated time | ||
| Timestep | ||
| Synaptic delay | ||
| Synaptic time constant | ||
| Decay of the STDP eligibility trace of presynaptic spikes | ||
| Decay of the STDP eligibility trace of postsynaptic spikes | ||
| STDP strengthening amplitude | ||
| STDP weakening amplitude | ||
| Maximum synaptic weight for STDP | ||
| Ring size | ||
| Random connections per cell |
We measured the resulting wallclock times on our hardware system ‘HWS2’ that we already used in subsection 3.7, using the parameter values from Table 7. Next, we determined the optimal runtime on this multi-core CPU system by selecting the best performance across the given ranges of MPI ranks and CPU threads, and subsequently compared the results for CoreNEURON, Arbor, and Arbor with STDP synapses (Fig. 9a-c). From this, we observe a tremendous speedup provided by Arbor, even if STDP dynamics are considered. Fig. 9d shows that the speedup is also maintained across network sizes, where both CoreNEURON and Arbor scale almost linearly. Furthermore, disentangling the setup and state propagation phases of the simulations, we find almost linear scaling for these particular measures as well (Fig. 9f). Exact runtime values, also for different dendritic tree depths, are provided in Table 8. Note that Arbor’s performance may be increased even further by appropriately adjusting the cpu_group_size setting for each particular case (here, all results are for cpu_group_size=1).
Regarding the memory use, we also observe a much better efficiency of Arbor as compared to CoreNEURON, while optimal results are achieved for the lowest rank and thread numbers (Supplementary Fig. S14). Again both simulators scale almost linearly (see Supplementary Table S1 for more details).
When additionally utilizing the existing GPU in the HWS2 setup described above, further speed gains can be achieved. In general, due to their large numbers of cores, it is expected that GPUs will accelerate simulations of large networks, which is supported by our results. Comparing Fig. 9f and Fig. 9g, we see that for large networks (here: 32768 neurons), especially if they also feature plasticity dynamics, the GPU brings benefits for both setup and propagation phase (see dark gray bars in Fig. 9g; also cf. Tables 8 and 9). However, for smaller networks of 1024 neurons, using the GPU can even slow down the simulation. For the exact runtime values with GPU, including such for different dendritic tree depths, see Table 9. Note that for CoreNEURON, none of the simulations finished because in all cases the compute system ran out of memory.
| Number of cells | Tree depth | Number of synapses | Number of compartments | Runtime of CoreNEURON (s) | Runtime of Arbor (s) | Runtime of Arbor with STDP (s) |
| 1024 | 0 | 1.03M | 1024 | 2.83 (0.75+2.07) | 0.23 (0.04+0.19) | 0.73 (0.05+0.68) |
| 2 | 1.03M | 46116 | 2.81 (0.78+2.04) | 0.32 (0.04+0.28) | 0.84 (0.05+0.79) | |
| 10 | 1.03M | 1672700 | 6.23 (0.65+5.58) | 2.67 (0.19+2.48) | 3.22 (0.21+3.01) | |
| 16384 | 0 | 16.41M | 16384 | 51.49 (5.28+46.21) | 3.24 (0.37+2.86) | 11.36 (0.75+10.61) |
| 2 | 16.41M | 737408 | 53.52 (5.32+48.20) | 4.64 (0.41+4.23) | 13.18 (0.81+12.37) | |
| 10 | 16.41M | 26616896 | 163.11 (7.47+155.64) | 43.07 (2.35+40.72) | 52.48 (2.58+49.90) | |
| 32768 | 0 | 32.81M | 32768 | 103.61 (9.57+94.04) | 6.11 (0.80+5.31) | 22.49 (1.48+21.00) |
| 2 | 32.81M | 1475268 | 106.17 (9.90+96.27) | 8.68 (0.82+7.86) | 26.16 (1.72+24.45) | |
| 10 | 32.81M | 52957704 | d.n.f. | 85.75 (4.65+81.09) | 104.66 (5.05+99.60) |
| Number of cells | Tree depth | Number of synapses | Number of compartments | Runtime of Arbor (s) | Runtime of Arbor with STDP (s) |
| 1024 | 0 | 1.03M | 1024 | 0.32 (0.02+0.30) | 0.81 (0.03+0.78) |
| 2 | 1.03M | 46116 | 0.44 (0.02+0.41) | 1.03 (0.03+1.00) | |
| 10 | 1.03M | 1672700 | 3.13 (0.14+2.99) | 3.65 (0.15+3.50) | |
| 16384 | 0 | 16.41M | 16384 | 3.10 (0.29+2.81) | 10.94 (0.50+10.44) |
| 2 | 16.41M | 737408 | 4.46 (0.32+4.15) | 12.60 (0.56+12.04) | |
| 10 | 16.41M | 26616896 | 43.51 (3.07+40.45) | 52.38 (3.33+49.05) | |
| 32768 | 0 | 32.81M | 32768 | 5.85 (0.55+5.30) | 21.51 (1.05+20.47) |
| 2 | 32.81M | 1475268 | 8.48 (0.70+7.78) | 24.84 (1.13+23.71) | |
| 10 | 32.81M | 52957704 | 86.62 (6.18+80.44) | 104.46 (6.71+97.75) |
By choosing different values for , , and , the workload can be adapted to investigate the network behavior of different real-world models. Considering different cell models further allows the emulation of their computational workload in network models. The relation of both network and cell parameters can then provide a holistic image of the performance of varying network models. In addition to the simple-branchy cell with Hodgkin-Huxley dynamics that we considered above, Arbor’s busyring implementation features another pre-configured cell model called complex. This model comprises a set of eight channel types on the soma, including kinetic schemes and calcium concentration models, as well as five channel types on the remaining morphology. In this way, the model approximates a cell from the mouse visual cortex [81]. A runtime comparison between complex and simple-branchy is shown in Supplementary Table S2.
The scaling behavior of Arbor enables the simulation of even larger and more complex network models than the ones we considered so far in this study. As an outlook, we show in Supplementary Fig. S15 first results of strong scaling on state-of-the-art supercomputing systems at JSC (the preview system JEDI and the current flagship system JUWELS) with an extremely large network of cells of the complex type (nevertheless, without plasticity dynamics). The parameters for both systems are given in Supplementary Table S3. As the NVIDIA H100 GPU model used in JEDI offers roughly twice the memory bandwidth and more than twice the floating-point performance of the A100 of JUWELS, we compare one JEDI node to two nodes of JUWELS. Within the strong scaling range over an eightfold increase of nodes in the simulation, we observe a scaling efficiency of well over 80%. The scaling efficiency is defined by
where is the wallclock time measured in the strong scaling series with nodes and reference .
4 Availability and Future Directions
4.1 Discussion and outlook